The product of twice a number and four is the same as the difference of seven times the number and 9/7. find the number.
Answer by checkley77(12844)

2x*4=7x-9/7
8x-7x=-9/7
x=-9/7 ans.
Proof:
2*-9/7*4=7*-9/7-9/7
-18/7*4=-63/7-9/7
-72/7=-72/7